OpenStudy (anonymous):

Can someone walk me through the steps to make sure that I am doing the right thing? Eliminate the parameter in the parametric equations x = (1/4)t and y = t2 and write the rectangular equation whose graph represents the curve.

OpenStudy (anonymous):

I think it is y=16x^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So that is the right answer?

OpenStudy (phi):

yes. x = (1/4) t so t = 4x (multiply both sides by 4) and t^2 = 16x^2 (square both sides) they say y = t^2 to y = 16 x^2

OpenStudy (anonymous):

Now the plus or minus from the square root goes away at the end, right?

OpenStudy (anonymous):

I doesn't matter, correct?

OpenStudy (phi):

what square root ? you start with x = (1/4)t and y = t^2 if you do as posted above, you never take the square root.

OpenStudy (anonymous):

Oh. Nevermind... Lol.

OpenStudy (anonymous):

Thanks. I looked at the wrong problem.

OpenStudy (phi):

you could do y = t^2 sqrt(y)= t x= (1/4)*sqrt(y) and you probably should take just the positive root

OpenStudy (anonymous):

Yeah, that's what I did... Ah! It's been a long week..., but thanks.